Diffusion Morning Report 03/07/11

Ted Berg tberg at stanford.edu
Mon Mar 7 06:25:33 PST 2011


Good Morning All,
     Looking Ok this morning . A few issues: Epi is down for chamber 
clean after Dome was coated, Tylan FGA is down for Temp issues, Tylan 1 
is down for Power fail on MFS. TEOS pump and Thermco LTO pumps out for 
repair.mco 2 no TLC flow. No other major issues at this time.

					
	Date: 	3/7/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*down* 	*power faIL* 	** 	** 	*3*
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*pump bad-out for repair* 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*15.53/30* 	** 	**
Tylan 10-Nitride 	*yellow* 	*door close issue* 	*7.89/30* 	** 	**
Tylan 11-LTO 	*up* 	*a bit thin* 	*20.26/30* 	** 	**
Tylan 12-FGA 	*down* 	*temp problems* 	** 	** 	*1*
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*down* 	*bad pump-out for repair* 	** 	** 	**
AMAT Epi 	*down* 	*chamber coated* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor going out for repair* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line 224 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110307/105e7640/attachment.html>


More information about the diffusion mailing list